OpenStudy (anonymous):

michele Given the parent functions f(x) = 3x + 2 and g(x) = 5^x –10, what is g(x) –f(x)?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute the subtraction: \[\Large {5^x} - 10 - \left( {3x + 2} \right) = ...?\]

OpenStudy (anonymous):

i don't know michele how to

OpenStudy (michele_laino):

hint: \[\Large {5^x} - 10 - \left( {3x + 2} \right) = {5^x} - 10 - 3x - 2 = ...?\]

OpenStudy (anonymous):

5^x-3x-10-2 ?

OpenStudy (michele_laino):

yes! and -10-2=-12

OpenStudy (anonymous):

-12=-12

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

g(x) –f(x) = 5^x –12 –3x this answer check if right?

OpenStudy (michele_laino):

yes! \[\Large {5^x} - 10 - \left( {3x + 2} \right) = {5^x} - 3x - 12\] that's right!

OpenStudy (anonymous):

good michele! thank

OpenStudy (michele_laino):

thanks!